OpenStudy (anonymous):

Solve this inequality: 4x+3=9 I'll need two points to graph this on a line. If someone can guide me through the steps to solve this, I'd appreciate it :)

OpenStudy (karatechopper):

this is an equation..

OpenStudy (karatechopper):

is there supposed to be some special sign for it?

OpenStudy (anonymous):

l4x+3l=9 is the original equation. It says to "Solve each open sentence. Then graph the solution set"

OpenStudy (mertsj):

|4x+3| = 9 means that 4x+3 is exactly 9 units from 0 Therefore 4x+3 = 9 or 4x+3=-9 Solve each of those open sentences

OpenStudy (karatechopper):

is that understood from what mert has said? or do u need more coaching

OpenStudy (anonymous):

Yep, I got it. However, how do I graph 3/2 on a graph? I'm talking about the graph that goes horizontally. Thanks :)

OpenStudy (mertsj):

You are talking about a number line graph. And 3/2 is 1 and 1/2 or 1.5

OpenStudy (mertsj):